# Med-R3: Enhancing Medical Retrieval-Augmented Reasoning of LLMs via Progressive Reinforcement Learning

### v3

@@ Appendix D Case Studies
Figure 8: Case study of Med-R3.
Appendix E Prompts
@@
Prompt: Content Suitability Judgment Please judge whether the following multiple-choice question is suitable for conversion into an open-ended question. The question to be converted must meet the following conditions: 1. After removing the options, the question itself remains valid, and the answer is unique and correct. 2. There is no ambiguity in the question and answer. 3. The question must have a unique optimal answer, not a range or a vague value. 4. Questions with negative options, such as selecting the option that does not meet the conditions or the “least likely” option, are not suitable. 5. In other cases, please use your logical judgment to make a decision. # Question
@@
Prompt: Reasoning Complexity Filtering You are an expert in the medical field. You will be given a question, a student’s answer, the correct and unique answer to the question, and other misleading options. Please compare the student’s answer with the correct answer and analyze whether the student’s answer is correct. In the student’s answer, the <think></think> tag wraps the thinking process, and the <answer></answer> tag wraps the final answer. During the model’s reasoning process, uncertain parts are encapsulated using the <search>search query</search> tags to facilitate subsequent information retrieval. Once relevant content is retrieved, the model continues its reasoning based on the retrieved information. It is important to note that the content within <document> search results </document> represents externally retrieved information and is not generated by the model itself. Therefore, this content should not be used to assess the model’s logical reasoning ability. Instead, it should be used solely to evaluate the model’s capability to process and integrate external information. Please judge the student’s thinking process and answer based on the correct answer, rate it from 1 to 5 points, and explain the reason. 5 Point Answer Criteria: 1. The thinking process is logical and seamless, and the reasoning process is specific and clear. 2. The final answer is consistent with the correct answer, allowing synonyms, abbreviations, etc. of the correct answer, but cannot contain incorrect options. 3-4 Point Answer Criteria: 1. The thinking process is reasonable. 2. The final answer is consistent with the correct answer, allowing some supplements, as long as they do not conflict with the correct answer, and the correct answer is the main one, not other misleading options. 1-2 Point Answer Criteria: 1. The thinking process is not clear. 2. The final answer is inconsistent with the correct answer or contains incorrect options. 3. Contains garbled characters, format errors, disorder, and irrelevant information. # Question
@@
Definitions: 1. entity1, relationship, entity2: Clearly extract entities or relationships from the natural language (e.g., “patient”, “has symptom”, “fever”). 2. if_retrieval: A boolean value (1/0) indicating whether the entity or relationship is retrieved from the search results within <document></document>. Processing rules: 1. Overall: Extract only entities and relationships related to solving medical problems, ignoring irrelevant background information. The knowledge graph should be concise and clear, avoiding redundancy, and based on facts, avoiding subjective speculation. 2. Entity Extraction: Must be specific medical concepts (diseases, symptoms, drugs, etc.), excluding vague descriptions (such as “some conditions” or “related factors”). Synonymous expressions should be merged (e.g., “myocardial infarction” instead of “heart attack”). 3. Relationship Definition: Use verb phrases (cause, inhibit, lead to, accompany, etc.) and must have a clear directionality (A→B or B→A should conform to medical logic). Negative relationships should be explicitly marked (e.g., “does not cause”, “rule out”). 4. Retrieval Marking Determination: If an entity or relationship is directly from the <document> search results </document>, mark “if_retrieval=1”; otherwise, “if_retrieval=0”. The search keywords themselves should not be considered as entities or relationships. 5. Special Handling: Retain hypothetical relationships in the reasoning process (marked as [hypothesis]), convert time relationships to medical temporal expressions (acute/chronic/ongoing period, etc.), and quantify probabilistic conclusions as (high/medium/low) risk levels.

### v4

@@ C.5. Environment and Hardware Configurations
Table 9. Meta-evaluation results across different evaluation strategies, including the exact match judge and the proprietary model DeepSeek-V3’s judge. We sample 30 instances per medical dataset.
Backbone Model Evaluation Strategy MedQA-USMLE MedQA-MCMLE MedMCQA RareArena-RDC RareArena-RDS MMLU-Med NEJMQA MedXpertQA Avg.
In-Domain (ID) Out-of-Domain (OOD)
Qwen2.5-7B Exact Match 0.80 (24/30) 0.70 (21/30) 0.73 (22/30) 0.30 (9/30) 0.37 (11/30) 0.60 (18/30) 0.53 (16/30) 0.43 (13/30) 0.56
LLM-as-Judge 1.00 (30/30) 0.97 (29/30) 1.00 (30/30) 0.97 (29/30) 1.00 (30/30) 0.93 (28/30) 0.93 (28/30) 0.97 (29/30) 0.97
LLaMA3.1-8B -Instruct Exact Match 0.70 (21/30) 0.60 (18/30) 0.67 (20/30) 0.40 (12/30) 0.33 (10/30) 0.70 (21/30) 0.50 (15/30) 0.53 (16/30) 0.55
LLM-as-Judge 1.00 (30/30) 1.00 (30/30) 1.00 (30/30) 0.93 (28/30) 0.97 (29/30) 1.00 (30/30) 0.90 (27/30) 0.93 (28/30) 0.97
Qwen2.5-14B Exact Match 0.83 (25/30) 0.67 (20/30) 0.76 (23/30) 0.60 (18/30) 0.57 (17/30) 0.67 (20/30) 0.80 (24/30) 0.63 (19/30) 0.69
LLM-as-Judge 1.00 (30/30) 1.00 (30/30) 0.97 (29/30) 1.00 (30/30) 0.97 (29/30) 1.00 (30/30) 0.97 (29/30) 0.90 (27/30) 0.98
Table 10. Failure case analysis with a held-out set of 300 questions from the MedQA-USMLE and MedMCQA datasets, where the model equipped with Med-R3 has produced incorrect answers. Here we employ Qwen2.5-14B for experimental conduction.
Category Sub-Category Failure Case
Component Content
Retrieval Failures (42%) Irrelevant Document Retrieval (28%) Question A 65-year-old man presents with “sudden-onset chest pain and shortness of breath.” Vital signs: BP 9060 mmHg, HR 120 bpm. Physical exam reveals bilateral pulmonary crackles. ECG shows ST-segment elevation.
Model Response The model retrieved a document discussing “gastroesophageal reflux disease as a cause of non-cardiac chest pain” and recommended upper endoscopy.
Error Analysis The term “chest pain” without clear cardiac context led the retriever to off-topic documents.
High-Quality Evidence Ignored (14%) Question Which medication is most likely to cause hypoglycemia in a patient with type 2 diabetes?
Model Response The model cited a 2020 observational study (Level IV evidence) claiming “rosiglitazone is associated with hypoglycemia”, while overlooking a 2023 large-scale RCT (Level I evidence) that clearly identified “glimepiride” as having the highest hypoglycemia risk.
Error Analysis The model prioritized recency over methodological rigor.
Reasoning Failures (58%) Misinterpretation of Evidence (33%) Question A literature excerpt states: “Long-term NSAID use is associated with an increased risk of gastric ulcers.” A patient with osteoarthritis has been taking ibuprofen chronically and now presents with melena.
Model Response The model concluded that “The melena is caused by ibuprofen-induced gastric ulcer; discontinue NSAIDs and switch to a COX-2 inhibitor.
Error Analysis The source only reported an association, not causation, and the model ignored other etiologies.
Incomplete Reasoning Chains (25%) Question A 40-year-old woman presents with fever, arthralgia, and rash. Lab findings: positive ANA (1:160), low complement C3.
Model Response The model directly diagnosed “systemic lupus erythematosus (SLE)” without considering alternative diagnoses.
Error Analysis The model failed to evaluate drug-induced lupus or mixed connective tissue disease.
C.6. Further Analysis on Proprietary Language Models
In our main experiment, we employ DeepSeek-V3 and DeepSeek-R1 as the proprietary language models for the following roles:
  1. (1)
    Training Data Construction: As descirbed in Section 2.1 and Figure 2, we employed the proprietary language models to first select the data that are suitable for RL from the original datasets, and then construct reference reasoning trajectories for the verification of reasoning process.
  2. (2)
    Evaluation: During the RL process, the proprietary language model DeepSeek-V3 serves as the judge of several reward components, e.g., entity/relation coverage via knowledge graph extraction, evidence quality (EBM level) scoring, etc. It is also adopted in the LLM-as-Judge paradigm to evaluate the correctness of models’ generated responses across medical benchmarks.
In this section, we present an in-depth analysis of the use of proprietary language models in our experimental pipeline, encompassing a systematic summary of their strengths, and human expert assessment of the judgments produced by these models.
C.6.1. Advantages of Proprietary Language Models
We employ the proprietary language models in our Med-R3 pipeline for the following reasons:
Reliability. LLM-as-Judge has become a standard evaluation approach for performance judgment, as human evaluation is costly and unscalable for open-ended tasks. This paradigm is now widely adopted in major LLM benchmarks where rule-based metrics fail to capture semantic correctness (Zheng et al., 2023; Li et al., 2025). Specifically, this strategy outperforms traditional metrics on complex medical reasoning tasks, since medical problems exhibit key characteristics that challenge conventional evaluation:
  • The answers are often expressed in diverse phrasings (no fixed template).
  • Correct responses may require justifying diagnoses or ruling out differential conditions.
  • There are multiple semantically equivalent formulations that can all be valid.
In such settings, exact-match against gold keys severely underestimates model capability (e.g., penalizing “myocardial infarction” as incorrect simply because the reference says “acute myocardial infarction”). Human evaluation, while ideal, is prohibitively expensive at scale and still subject to inter-annotator disagreement, even among clinicians. In contrast, LLM-as-Judge enables semantic equivalence judgment and logical coherence assessment, aligning better with the open-ended nature of clinical reasoning. Moreover, our evaluation process mitigates the risk of judge bias by enforcing a fixed and well-defined output format.
Robustness. Moreover, during the training data construction phase, particularly in the data filtering process based on reasoning complexity and in generating reference reasoning trajectories for medical questions to support subsequent RL verification, we first used DeepSeek-R1 to generate multiple candidate answers, and then employed DeepSeek-V3 to evaluate the quality of these sequences generated by DeepSeek-R1, thereby enhancing fault tolerance.
Reproducibility. We have also provided the prompts used during data construction and performance judgment, including those guiding content validity assessment, reasoning difficulty–based sample selection, question rewriting, reward verification in the RL process, and answer correctness evaluation, as detailed in Appendix G.
C.6.2. Meta-Evaluation of Proprietary Language Models
Here we have conducted a small-scale validation study to evaluate the quality of proprietary language model in comparison with human-annotation (we employ DeepSeek-V3 in our main experiment) during the data curation and RL process.
  • Entity/Relation Extraction Accuracy (Equation 1): We randomly sampled 200 reasoning steps from the training trajectories and asked three board-certified physicians to annotate entities and relations according to our schema (e.g., “A causes B”, “A inhibits B”). Compared to human annotations, DeepSeek-V3’s extractions achieve:
    • Entity: 92.3% (F1 score)
    • Relation (directional): 86.7% (F1 score)
    The inter-annotator agreement (Fleiss’ k (Fleiss, 1971)) was 0.81, indicating substantial agreement.
  • EBM Level Agreement (Equation 10): For 150 retrieved documents from the external knowledge corpus, we compared DeepSeek-V3’s EBM level assignments (Level I–VI) against consensus labels from two independent clinical researchers. The weighted Cohen’s k (Cohen, 1968) was 0.74, reflecting strong agreement, especially for high-quality evidence (Level I–II), where the precision exceeded 89%.
  • Answer Correctness Judgment: We have also held the human meta-evaluation study to assess the judge quality of different evaluation strategies, including exact-match and LLM-as-Judge. Specifically, we sample 30 instances from each medical dataset, collect the generation results from Qwen2.5-7B, LLaMA3.1-8B-Instruct and Qwen2.5-14B, and then report the judgment correctness of exact-match and LLM-as-Judge from the proprietary model DeepSeek-V3 (employed in our main experiment). As depicted in Table 9, the experimental results demonstrate that the proprietary model evaluator achieves approximately 97% agreement with human experts, whereas the exact-match judge attains only about 60% alignment on average. On benchmarks related to real-world clinical scenarios, such as RareArena, MedXpertQA, and NEJMQA, the exact-match judge exhibits even lower alignment with human experts. It is primarily because questions in these datasets are often open-ended or admit multiple clinically valid answers, whereas exact-match evaluation only considers literal string identity and fails to account for semantic equivalence or acceptable paraphrasing in medical reasoning. This further underscores the reliability of LLM-as-Judge.
@@ Appendix D Case Studies
Figure 8. Case study of Med-R3.
We have provided case studies in Figure 8. Below we describe the core improvements introduced by our method:
  • Proactive Retrieval Triggering: The model doesn’t just retrieve when prompted—it identifies knowledge gaps (e.g., “Is asbestos exposure more strongly linked to mesothelioma or lung cancer?”) and initiates search autonomously.
  • Evidence-Based Decision Making: Instead of relying on memorized associations, it integrates structured evidence (e.g., SMR values, study populations) into its reasoning chain.
  • Multi-Factor Risk Integration: It combines occupational history, smoking status, symptom progression, and epidemiological data to reach a balanced conclusion.
Appendix E Failure Case Analysis
We have expanded our analysis of failure cases using a held-out set of 300 questions from the MedQA-USMLE and MedMCQA datasets where Med-R3 has produced incorrect answers (the backbone model we employ here is Qwen2.5-14B, i.e., Qwen2.5-14B + Med-R3). Two board-certified physicians independently categorized errors into retrieval and reasoning types, with a Fleiss’ k (Fleiss, 1971) of 0.79. Our key observations are outlined in Table 10.
The failure cases can be categorized into the following types:
  • Retrieval Failures (42%)
    • Irrelevant Document Retrieval (28%): Sometimes the ambiguous symptom descriptions lead to off-topic retrieved documents.
    • High-Quality Evidence Ignored (14%): The model sometimes prioritizes recent but lower-evidence papers over systematic reviews or RCTs.
  • Reasoning Failures (58%)
    • Misinterpretation of Evidence (33%): Although the model has retrieved correct documents, it confuses correlation with causation.
    • Incomplete Reasoning Chains (25%): The model gives premature conclusion without integrating multi-step clinical logic.
Appendix F Ethical Considerations
Appendix G Prompts
@@
Prompt: Content Suitability Judgment Please judge whether the following multiple-choice question is suitable for conversion into an open-ended question. The question to be converted must meet the following conditions: (1) After removing the options, the question itself remains valid, and the answer is unique and correct. (2) There is no ambiguity in the question and answer. (3) The question must have a unique optimal answer, not a range or a vague value. (4) Questions with negative options, such as selecting the option that does not meet the conditions or the “least likely” option, are not suitable. (5) In other cases, please use your logical judgment to make a decision. # Question
@@
Prompt: Reasoning Complexity Filtering You are an expert in the medical field. You will be given a question, a student’s answer, the correct and unique answer to the question, and other misleading options. Please compare the student’s answer with the correct answer and analyze whether the student’s answer is correct. In the student’s answer, the <think></think> tag wraps the thinking process, and the <answer></answer> tag wraps the final answer. During the model’s reasoning process, uncertain parts are encapsulated using the <search>search query</search> tags to facilitate subsequent information retrieval. Once relevant content is retrieved, the model continues its reasoning based on the retrieved information. It is important to note that the content within <document> search results </document> represents externally retrieved information and is not generated by the model itself. Therefore, this content should not be used to assess the model’s logical reasoning ability. Instead, it should be used solely to evaluate the model’s capability to process and integrate external information. Please judge the student’s thinking process and answer based on the correct answer, rate it from 1 to 5 points, and explain the reason. 5 Point Answer Criteria: (1) The thinking process is logical and seamless, and the reasoning process is specific and clear. (2) The final answer is consistent with the correct answer, allowing synonyms, abbreviations, etc. of the correct answer, but cannot contain incorrect options. 3-4 Point Answer Criteria: (1) The thinking process is reasonable. (2) The final answer is consistent with the correct answer, allowing some supplements, as long as they do not conflict with the correct answer, and the correct answer is the main one, not other misleading options. 1-2 Point Answer Criteria: (1) The thinking process is not clear. (2) The final answer is inconsistent with the correct answer or contains incorrect options. (3) Contains garbled characters, format errors, disorder, and irrelevant information. # Question
@@
Definitions: (1) entity1, relationship, entity2: Clearly extract entities or relationships from the natural language (e.g., “patient”, “has symptom”, “fever”). (2) if_retrieval: A boolean value (1/0) indicating whether the entity or relationship is retrieved from the search results within <document></document>. Processing rules: (1) Overall: Extract only entities and relationships related to solving medical problems, ignoring irrelevant background information. The knowledge graph should be concise and clear, avoiding redundancy, and based on facts, avoiding subjective speculation. (2) Entity Extraction: Must be specific medical concepts (diseases, symptoms, drugs, etc.), excluding vague descriptions (such as “some conditions” or “related factors”). Synonymous expressions should be merged (e.g., “myocardial infarction” instead of “heart attack”). (3) Relationship Definition: Use verb phrases (cause, inhibit, lead to, accompany, etc.) and must have a clear directionality (A→B or B→A should conform to medical logic). Negative relationships should be explicitly marked (e.g., “does not cause”, “rule out”). (4) Retrieval Marking Determination: If an entity or relationship is directly from the <document> search results </document>, mark “if_retrieval=1”; otherwise, “if_retrieval=0”. The search keywords themselves should not be considered as entities or relationships. (5) Special Handling: Retain hypothetical relationships in the reasoning process (marked as [hypothesis]), convert time relationships to medical temporal expressions (acute/chronic/ongoing period, etc.), and quantify probabilistic conclusions as (high/medium/low) risk levels.